A tank can be filled in 2 hours. The same tank can be emptied in 6 hours. If the faucet is on and the drain is left open, when will the tank overflow

Answer by nerdybill(7384)

A tank can be filled in 2 hours. The same tank can be emptied in 6 hours. If the faucet is on and the drain is left open, when will the tank overflow
.
Let x = time (hours) when tank overflows
then
x(1/2 - 1/6) = 1
multiplying both sides by 6:
x(3 - 1) = 6
x(2) = 6
x = 6/2
x = 3 hours